OpenStudy (anonymous):

how many possible combinations can there be with the numbers 6,6,4,4,1?

OpenStudy (anonymous):

if they were all different you would have 6! combinations, but since you cannot tell the 6s apart, nor the twos, it is \[\frac{6!}{2\times2}\]

OpenStudy (anonymous):

"6,6,4,4,1" are 5 numbers isn't ?

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

yeah i guess it is isn't it!

OpenStudy (anonymous):

so if it was 42333 what would it be? whats the general formula?

OpenStudy (anonymous):

what is an additional number between friends? ok i was wrong, maybe try \[\frac{5!}{2\times 2}\] if you want to actaully get the correct answer

OpenStudy (anonymous):

haha :D

OpenStudy (anonymous):

\[\frac{5\times 4\times 3\times 2}{2\times 2}=5\times 3\times 2=30\] (did i get that right?)

OpenStudy (anonymous):

looks correct, but what is the general formula?

OpenStudy (anonymous):

\[\frac{5!}{3!}\] for the second one

OpenStudy (anonymous):

so that otheer one is really 2!x2! in the denominator?

OpenStudy (anonymous):

yes, if you want to think of it that way. the number of ways you can permute 2 things is 2! = 2

OpenStudy (anonymous):

awesome thanks so much!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

sorry about mis-counting

OpenStudy (anonymous):

its all good... all worked out!